Tech Staff Meeting Notes October 31, 2016
The top priority is always repair of down stations.
NM problem stations
ET problem stations
Duty Table updated June 2014
This was an abbreviated meeting prior to SIS training. The primary topics were the GX620 motherboard Capacitor review and Jim's priority list. Steve will construct list of all GX620's and indicate the results of a visual inspection of every machine, both in house and deployed.
Presentations from the NetOpsVIII workshop are online at http://www.ceri.memphis.edu/people/mwithers/NetOpsVIII/.
Pramitol is condisered a Hazardous Chemical. This email chain from Al at ES&H provides permission to purchase it.
Some GX620 machines have bad capacitors on the motherboard. Dave found a source for motherboards (sans cpu and ram) and powersupplies. A new one is operational in nmadworm. He'll get six to ten more.
NMEM is now an obsidian.
2016 Routine Service is complete for all stations that need it.
EPRM -dropping out
SJBM - dropping out
BOAR - "
CPAR - "
LVAR - "
RVAR - "
TYAR - "
LRAR - "
CCAR.HNE - drifting
FCAR.HN* see image
LCAR.HN* see image
WYBT was vandalized for the second year in a row. Chris got it going temporarily. Need to revisit and harden.
MORT is closed. Property owner plans to continue farming behind the levee and keep the road open. Will reinstall flood resistant MORT2 at a nicer spot.
There was an email discussion on the history of OXF which led to digging up the MOU from 1993. Mitch checked with Dave Wilson at ASL and it may be possible to replace the MOU, which is endorsed at a fairly high level within USGS, with a "station agreement" which is less formal.
Steve asked about replacing all node disks with solid state disk. Funds are available. Steve will get an estimate.
All nodes are due a firmware upgrade that will sufficiently change things such that the the initial upgrade and configure should be done at CERI, then swap hardware at the node.
John checked other networks plans to replace Windows digiziting boxes and found none. NC has migrated to the NI X-series boards and adsend does work with it though its still on MS Windows. John will try to find out which specific board and OS is used.
DBB: Move the DBB link at CERI from the west end of 3892 to the east side of 3904 and eliminate the thin-wire. Tower is up to test the waters.
Uverse is likely available at the lnxt node. Steve will investigate to see if its cost effective to upgrade from dsl. We don't need the band width but service for dsl will likely continue to deteriorate.
Flooding in the expansion space has become a problem. A Workorder was submitted and P-plant engineers were on-site to assess the problem. The most likely solution will be a retention wall and removal of material to make the foundation above grave.
We discovered that most of the ISIS4 packages have a wire reversal that inverts the orientations. All but two ISIS4 stations are impacted (RCGA and DSTN). John will visit and repair each station. John will also create a spreadsheet that lists each station, when the ISIS4 was intalled and when it was repaired so that Mitch can correct the orientations in the metadata.
Replacement station for GTTN and RBNC being evaluated and permissioned for Southern Highlands Science Learning Center at Purchase Knob, a part of the Great Smoky Mountains National Park. Noise study picked up compressor noise from a nearby air quality station. Need to add isolation mounts and test farther away. We had to return the compact so will need to use a 6TD instead. Cell service is probably not available so a non-real-time 6TD is needed.
ASTN.HN* see image
GRBT.EHE see image
Drifting sensor at FPAL was replaced. Its behaving normally in the test vault. Nanometrics has not yet provided a quote to upgrade the PA to QA.
GMG has had a power upgrade, new batteries, new solar contoller, and better surge suppression. And it still went down on 1/18. James can get to the radio at the station which is on a separate power feed than the reftek. Jim will send a new BC10 and John will investigate.
Obsidian and fanless PC for CVVA is mocked up and ready for installation.
John's updated todo list .
Improve RCGA AC power and perform maintenance work on shed.
Move PMTN due to excessive noise and vandalism. The trees at Lone Mountian are too tall to get a signal out. Looking for an alternate site.
HPNC node is up. Need to fix Win7 to allow earthworm commands (e.g. status, restart etc) from the ssh command line. 9 channels(GFM, MGNC, and DSNC) are being exported over a cell modem temporarily until internet is established. Still waiting on completion of communcations to get on local internet. Its still on a cell modem; Randy Loop plans to install the internet modem next week. Hopefully after that we can retrieve the cell modem and get all channels to CERI in real-time.
Upgrade remaining High Peak Node Stations from L4-3D's to S-13's. At least one bad component in L4-3's at CCNC, TVNC, TRYN.
GFM2 is good to go. It will be a reftek with S-13's. Do we need to take student volunteers to help with the GFM decommission?
WMTN pit needs to be moved due to erosion.
CPRT comm reconfig. John says, "CPRT is going through the Senior Center at Corryton (cort). We removed everything from svfd an everything but what is on the tower at gttn. cort is indeed using dynamic addressing but data is flowing. Greg Hill the director at the senior center is supposed to be contacting his IT person to see what can be done. I left a message for Greg last week but still haven't heard back. I want to get Greg's IT guy in direct contact with either Steve or James so they can figure it all out."
Bob and Chris found a probable site for the new station and are working on a permit.
CVMS has intermittent connection problems with wireless. Its hardwired for now but the wireless needs to be replaced to preserve electical isolation between the building and vault.
All GPS stations are back up.
Z41A is online with AG. Vault leaks so did not install sm sensor. Need to fix the leak and install the sm vault inside the TA vault.
When the water leak at LCAR was repaired, was an accelerometer reinstalled?
Needs some clean up.
The FDL wiki is here: https://umwiki.memphis.edu/display/FDL/CERI+FDL
The seisnet group on umdrive is at http://umdrive.memphis.edu/mwithers/seisnet You will need your uuid and password.
The seisnet wiki is at https://umwiki.memphis.edu/display/CERITech/CERI+Tech You will need your uuid and password.